# Condensed-space methods for nonlinear programming on GPUs

### v1

@@ 5.2.2 Tuning the Golub & Greif strategy
In Figure 1 we depict the evolution of the number of Cg iterations and relative accuracy as we increase the parameter $\gamma$ from $10^{4}$ to $10^{8}$ in HyKKT.
On the algorithmic side, we observe that the higher the regularization $\gamma$, the faster the Cg algorithm: we decrease the total number of iterations spent in Cg by a factor of 10. However, we have to pay a price in term of accuracy: for $\gamma>10^{8}$ the solution returned by the linear solver is not accurate enough and the IPM algorithm has to proceed to additional primal-dual regularizations.
On the numerical side, the table in Figure 1 compares the time spent in the IPM solver on the CPU (using CHOLMOD) and on the GPU (using the solver cuDSS). Overall cuDSS is faster than CHOLMOD, leading to a 4x-8x speed-up in the total IPM solution time. We note also that the assembly of the condensed matrix $K_{\gamma}$ parallelizes well on the GPU, with a reduction in the assembly time from $\approx 8s$ on the CPU to $\approx 0.2s$ on the GPU.
CHOLMOD-$\mathrm{LDL^{T}}$ (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\gamma$ it cond. (s) Cg (s) linsol (s) IPM (s) # it cond. (s) Cg (s) linsol (s) IPM (s)
$10^{4}$ 96 8.15 463.86 536.83 575.06 96 0.17 113.27 114.52 124.00
$10^{5}$ 96 8.33 163.35 235.61 273.36 96 0.17 53.37 54.62 64.39
$10^{6}$ 96 8.22 68.69 139.86 177.24 96 0.17 14.53 15.78 25.39
$10^{7}$ 96 8.24 35.12 107.17 144.78 96 0.17 7.95 9.20 18.41
$10^{8}$ 96 7.89 21.68 93.85 131.33 96 0.17 5.36 6.62 15.90
Figure 1: Above: Decomposition of IPM solution time across (a) condensation time (cond.), (b) Cg time, (c) total time spent in the linear solver (linsol.) and (d) total time spent in IPM solver (IPM). Below: Impact of $\gamma$ on the total number of Cg iterations and the norm of the relative residual at each IPM iteration. The peak observed in the norm of the relative residual corresponds to the primal-dual regularization performed inside the IPM algorithm, applied when the matrix $K_{\gamma}$ is not positive definite.
@@ 5.2.3 Tuning the equality relaxation strategy
We now analyze the numerical performance of LiftedKKT (§3.2). The method solves the KKT system (13) using a direct solver. The parameter $\tau$ used in the equality relaxation (12) is set equal to the IPM tolerance $\varepsilon_{tol}$ (in practice, it does not make sense to set a parameter $\tau$ below IPM tolerance as the inequality constraints are satisfied only up to a tolerance ±$\pm\varepsilon_{tol}$ in IPM).
We compare in Table 2 the performance obtained by LiftedKKT as we decrease the IPM tolerance $\varepsilon_{tol}$. We display both the runtimes on the CPU (using CHOLMOD-$\mathrm{LDL^{T}}$) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$). The slacks associated with the relaxed equality constraints are converging to a value below $2\tau$, leading to highly ill-conditioned terms in the diagonal matrices $D_{s}$. As a consequence, the conditioning of the matrix $K_{\tau}$ in (13) can increase above $10^{18}$, leading to a nearly singular linear system. We observe cuDSS-$\mathrm{LDL^{T}}$ is more stable than CHOLMOD: the factorization succeeds, and the loss of accuracy caused by the ill-conditioning is tamed by the multiple Richardson iterations that reduces the relative accuracy in the residual down to an acceptable level. As a result, cuDSS can solve the problem to optimality in $\approx 20s$, a time comparable with HyKKT (see Figure 1).
CHOLMOD-$\mathrm{LDL^{T}}$ (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\varepsilon_{tol}$it time (s) #it time (s) accuracy
$10^{-4}$ 115 268.2 114 19.9 $1.2\times 10^{-2}$
$10^{-5}$ 210 777.8 113 30.4 $1.2\times 10^{-3}$
$10^{-6}$ 102 337.5 109 25.0 $1.2\times 10^{-4}$
$10^{-7}$ 108 352.9 104 20.1 $1.2\times 10^{-5}$
$10^{-8}$ 105 20.3 $1.2\times 10^{-6}$
Table 2: Performance of the equality-relaxation strategy as we decrease the IPM tolerance $\varepsilon_{tol}$. The table displays the wall time on the CPU (using CHOLMOD-$\mathrm{LDL^{T}}$) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$).
@@ 5.2.4 Breakdown of the time spent in one IPM iteration
We decompose the time spent in a single IPM iteration for LiftedKKT and HyKKT. As a reference running on the CPU, we show the time spent in the solver HSL MA27. We observe that HSL MA57 is slower than HSL MA27, as the OPF instances are super-sparse. Hence, the block elimination algorithm implemented in HSL MA57 is not beneficial there  [^2] Personal communication with Iain Duff..
When solving the KKT system, the time can be decomposed into: (1) assembling the KKT system, (2) factorizing the KKT system, and (3) computing the descent direction with triangular solves. As depicted in Figure 2, we observe that constructing the KKT system represents only a fraction of the computation time, compared to the factorization and the triangular solves. Using cuDSS-$\mathrm{LDL^{T}}$, we observe speedups of 30x and 15x in the factorization compared to MA27 and CHOLMOD running on the CPU. Once the KKT system is factorized, computing the descent direction with LiftedKKT is faster than with HyKKT (0.04s compared to 0.13s) as HyKKT has to run a Cg algorithm to solve the Schur complement system (11), leading to additional backsolves in the linear solver.
build (s) factorize (s) backsolve (s) accuracy
HSL MA27 $3.15\times 10^{-2}$ $1.22\times 10^{-0}$ $3.58\times 10^{-1}$ $5.52\times 10^{-7}$
LiftedKKT (CPU) $8.71\times 10^{-2}$ $6.08\times 10^{-1}$ $2.32\times 10^{-1}$ $3.73\times 10^{-9}$
HyKKT (CPU) $7.97\times 10^{-2}$ $6.02\times 10^{-1}$ $7.30\times 10^{-1}$ $3.38\times 10^{-3}$
LiftedKKT (CUDA) $2.09\times 10^{-3}$ $4.37\times 10^{-2}$ $3.53\times 10^{-2}$ $4.86\times 10^{-9}$
HyKKT (CUDA) $1.86\times 10^{-3}$ $3.38\times 10^{-2}$ $1.35\times 10^{-1}$ $3.91\times 10^{-3}$
Figure 2: Breakdown of the time spent in one IPM iteration for different linear solvers, when solving 78484epigrids (A30 GPU)
@@ 5.3 Benchmark on OPF instances
We run a benchmark on difficult OPF instances taken from the PGLIB benchmark babaeinejadsarookolaee2019power . We compare LiftedKKT and HyKKT with HSL MA27. The results are displayed in Table 3, for an IPM tolerance set to $10^{-6}$. Regarding HyKKT, we set $\gamma=10^{7}$ following the analysis in §5.2.2. The table displays the time spent in the initialization, the time spent in the linear solver and the total solving time. We complement the table with a Dolan & Moré performance profile dolan2002benchmarking displayed in Figure 3. Overall, the performance of HSL MA27 on the CPU is consistent with what was reported in babaeinejadsarookolaee2019power .
@@
On the GPU, LiftedKKT+cuDSS is faster than HyKKT+cuDSS on small and medium instances: indeed, the algorithm does not have to run a Cg algorithm at each IPM iteration, limiting the number of triangular solves performed at each iteration. Both LiftedKKT+cuDSS and HyKKT+cuDSS are significantly faster than HSL MA27. HyKKT+cuDSS is slower when solving 8387_pegase, as on this particular instance the parameter $\gamma$ is not set high enough to reduce the total number of Cg iterations, leading to a 4x slowdown compared to LiftedKKT+cuDSS. Nevertheless, the performance of HyKKT+cuDSS is better on the largest instances, with almost an 8x speed-up compared to the reference HSL MA27.
@@
The benchmark presented in Table 3 has been generated using a NVIDIA A100 GPUs (current selling price: \$10k). We have also compared the performance with cheaper GPUs: a NVIDIA A1000 (a laptop-based GPU, 4GB memory) and a NVIDIA A30 (24GB memory, price: \$5k). As a comparison, the selling price of the AMD EPYC 7443 processor we used for the benchmark on the CPU is \$1.2k. The results are displayed in Figure 4. We observe that the performance of the A30 and the A100 are relatively similar. The cheaper A1000 GPU is already faster than HSL MA27 running on the CPU, but is not able to solve the largest instance as it is running out of memory.
HSL MA27 LiftedKKT+cuDSS HyKKT+cuDSS
Case it init lin total it init lin total it init lin total
89_pegase 32 0.00 0.02 0.03 29 0.03 0.12 0.24 32 0.03 0.07 0.22
179_goc 45 0.00 0.03 0.05 39 0.03 0.19 0.35 45 0.03 0.07 0.25
500_goc 39 0.01 0.10 0.14 39 0.05 0.09 0.26 39 0.05 0.07 0.27
793_goc 35 0.01 0.12 0.18 57 0.06 0.27 0.52 35 0.05 0.10 0.30
1354_pegase 49 0.02 0.35 0.52 96 0.12 0.69 1.22 49 0.12 0.17 0.50
2000_goc 42 0.03 0.66 0.93 46 0.15 0.30 0.66 42 0.16 0.14 0.50
2312_goc 43 0.02 0.59 0.82 45 0.14 0.32 0.68 43 0.14 0.21 0.56
2742_goc 125 0.04 3.76 7.31 157 0.20 1.93 15.49 - - - -
2869_pegase 55 0.04 1.09 1.52 57 0.20 0.30 0.80 55 0.21 0.26 0.73
3022_goc 55 0.03 0.98 1.39 48 0.18 0.23 0.66 55 0.18 0.23 0.68
3970_goc 48 0.05 1.95 2.53 47 0.26 0.37 0.87 48 0.27 0.24 0.80
4020_goc 59 0.06 3.90 4.60 123 0.28 1.75 3.15 59 0.29 0.41 1.08
4601_goc 71 0.09 3.09 4.16 67 0.27 0.51 1.17 71 0.28 0.39 1.12
4619_goc 49 0.07 3.21 3.91 49 0.34 0.59 1.25 49 0.33 0.31 0.95
4837_goc 59 0.08 2.49 3.33 59 0.29 0.58 1.31 59 0.29 0.35 0.98
4917_goc 63 0.07 1.97 2.72 55 0.26 0.55 1.18 63 0.26 0.34 0.94
5658_epigrids 51 0.31 2.80 3.86 58 0.35 0.66 1.51 51 0.35 0.35 1.03
7336_epigrids 50 0.13 3.60 4.91 56 0.45 0.95 1.89 50 0.43 0.35 1.13
8387_pegase 74 0.14 5.31 7.62 82 0.59 0.79 2.30 75 0.58 7.66 8.84
9241_pegase 74 0.15 6.11 8.60 101 0.63 0.88 2.76 71 0.63 0.99 2.24
9591_goc 67 0.20 11.14 13.37 98 0.63 2.67 4.58 67 0.62 0.74 1.96
10000_goc 82 0.15 6.00 8.16 64 0.49 0.81 1.83 82 0.49 0.75 1.82
10192_epigrids 54 0.41 7.79 10.08 57 0.67 1.14 2.40 54 0.67 0.66 1.81
10480_goc 71 0.24 12.04 14.74 67 0.75 0.99 2.72 71 0.74 1.09 2.50
13659_pegase 63 0.45 7.21 10.14 75 0.83 1.05 2.96 62 0.84 0.93 2.47
19402_goc 69 0.63 31.71 36.92 73 1.42 2.28 5.38 69 1.44 1.93 4.31
20758_epigrids 51 0.63 14.27 18.21 53 1.34 1.05 3.57 51 1.35 1.55 3.51
30000_goc 183 0.65 63.02 75.95 - - - - 225 1.22 5.59 10.27
78484_epigrids 102 2.57 179.29 207.79 101 5.94 5.62 18.03 104 6.29 9.01 18.90
Table 3: OPF benchmark, solved with a tolerance tol=1e-6. (A100 GPU)
@@
Figure 3: Performance profile for the PGLIB OPF benchmark, solved with a tolerance tol=1e-6.
@@
Figure 4: Comparing the performance obtained with various GPUs on three different OPF instances.
5.4 Benchmark on COPS instances
We have observed in the previous section that both LiftedKKT and HyKKT outperforms HSL MA27 when running on the GPU. However, the OPF instances are specific nonlinear instances. For that reason, we complement our analysis by looking at the performance of LiftedKKT and HyKKT on large-scale COPS instances dolan2004benchmarking . We look at the performance we get on the COPS instances used in the Mittelmann benchmark mittelmann2002benchmark . To illustrate the heterogeneity of the COPS instances, we display in Figure 5 the sparsity pattern of the condensed matrices $K_{\gamma}$ (10) for one OPF instance and for multiple COPS instances. We observe that some instances (bearing) have a sparsity pattern similar to the OPF instance on the left, whereas some are fully dense (elec). On the opposite, the optimal control instances (marine, steering) are highly sparse and can be reordered efficiently using AMD ordering amestoy-david-duff-2004 .
The results of the COPS benchmark are displayed in Table 4. HSL MA57 gives better results than HSL MA27 for the COPS benchmark, and for that reason we have decided to replace HSL MA27 by HSL MA57. As expected, the results are different than on the OPF benchmark. We observe that LiftedKKT+cuDSS and HyKKT+cuDSS outperform HSL MA57 on the dense instance elec (20x speed-up) and bearing — an instance whose sparsity pattern is similar to the OPF. In the other instances, LiftedKKT+cuDSS and HyKKT+cuDSS on par with HSL MA57 and sometimes even slightly slower (rocket and pinene).
Figure 5: Sparsity patterns for one OPF instance and for various COPS problems. The first row displays the sparsity pattern of $K_{\gamma}$, after AMD reordering. The second row displays the sparsity pattern of $K_{\gamma}$ after Metis reordering.

### v2

@@ 6.1.2 Tuning the Golub & Greif strategy
We observe that the higher the regularization $\gamma$, the faster the Cg algorithm. However, we have to pay a price in terms of accuracy. In Figure 2, we depict the evolution of the number of Cg iterations and relative accuracy as we increase the parameter $\gamma$ from $10^{4}$ to $10^{8}$ in HyKKT. We observe that the number of CG iterations decreases by a factor of 10 as we increase $\gamma$, but the norm of the relative residual increases linearly as we increase $\gamma$. Thus, a trade-off exists between the time spent in Cg and the accuracy of the solution.
On the numerical side, the table in Figure 2 compares the time spent in the IPM solver on the CPU (using Pardiso ) and on the GPU (using the solver cuDSS). Overall, cuDSS is faster than Pardiso , leading to a 4x-8x speed-up in the total IPM solution time. We also note that the assembly of the condensed matrix $K_{\gamma}$ parallelizes well on the GPU, with a reduction in the assembly time from $\approx 8s$ on the CPU to $\approx 0.2s$ on the GPU.
Pardiso (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\gamma$ # it cond. (s) Cg (s) linsol (s) IPM (s) # it cond. (s) Cg (s) linsol (s) IPM (s)
$10^{4}$ 96 8.1 562.1 599.8 639.2 96 0.17 113.27 114.52 124.00
$10^{5}$ 96 8.5 212.0 249.9 290.0 96 0.17 53.37 54.62 64.39
$10^{6}$ 96 8.6 96.7 134.6 174.6 96 0.17 14.53 15.78 25.39
$10^{7}$ 96 8.6 52.9 91.0 130.5 96 0.17 7.95 9.20 18.41
$10^{8}$ 96 8.6 35.1 73.2 113.3 96 0.17 5.36 6.62 15.90
Figure 2: Above: Decomposition of IPM solution time across (a) condensation time (cond.), (b) Cg time, (c) total time spent in the linear solver (linsol.) and (d) total time spent in IPM solver (IPM). We compare the performance of cuDSS with those obtained by Pardiso running in parallel on the CPU using 8 threads. Below: Impact of $\gamma$ on the total number of Cg iterations and the norm of the relative residual at each IPM iteration. The peak observed in the norm of the relative residual corresponds to the primal-dual regularization performed inside the IPM algorithm, applied when the matrix $K_{\gamma}$ is not positive definite. The column for it displays the number of iterations, whereas the columns for cond., CG, linsol and IPM display the time spent in assembling the condensed KKT matrix, the time spent in the conjugate gradient, the time spent in the linear solver (including CG) and the total time spent in the IPM algorithm, respectively.
@@ 6.1.3 Tuning the equality relaxation strategy
We now analyze the numerical performance of LiftedKKT (§3.2). The method solves the KKT system (15) using a direct solver. The parameter $\tau$ used in the equality relaxation (14) is set equal to the IPM tolerance $\varepsilon_{tol}$ (in practice, it does not make sense to set a parameter $\tau$ below IPM tolerance as the inequality constraints are satisfied only up to a tolerance $\pm\varepsilon_{tol}$ in IPM).
We compare in Table 2 the performance obtained by LiftedKKT as we decrease the IPM tolerance $\varepsilon_{tol}$. We display both the runtimes on the CPU (using Pardiso with Cholesky) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$). The slacks associated with the relaxed equality constraints are converging to a value below $2\tau$, leading to highly ill-conditioned terms in the diagonal matrices $D_{s}$. As a consequence, the conditioning of the matrix $K_{\tau}$ in (15) can increase above $10^{18}$, leading to a nearly singular linear system. However, the factorization succeeds and the loss of accuracy caused by the ill-conditioning is tamed by the multiple Richardson iterations in the iterative refinement, reducing the relative accuracy in the residual down to an acceptable level. As a result, cuDSS can solve the problem to optimality in $\approx 20s$, a time comparable with HyKKT (see Figure 2). We observe that updating the parameter $\tau$ changes the total number of iterations in IPM: this is expected, as we solve a different problem (14) every time we update the regularization $\tau$.
Pardiso-$\mathrm{LDL^{T}}$ (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\varepsilon_{tol}$ #it time (s) #it time (s) accuracy
$10^{-4}$ 127 164.9 114 19.9 $1.2\times 10^{-2}$
$10^{-5}$ 129 168.1 113 30.4 $1.2\times 10^{-3}$
$10^{-6}$ 200 301.7 109 25.0 $1.2\times 10^{-4}$
$0^{-7} 11$0 159.2 104 20.1 $1.2\times 10^{-5}$
$0^{-8} 11$1 156.9 105 20.3 $1.2\times 10^{-6}$
Table 2: Performance of the equality-relaxation strategy as we decrease the IPM tolerance $\varepsilon_{tol}$. The table displays the wall time on the CPU (using Pardiso-$\mathrm{LDL^{T}}$) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$).
@@ 6.1.4 Breakdown of the time spent in one IPM iteration
We decompose the time spent in a single IPM iteration for LiftedKKT and HyKKT. As a reference running on the CPU, we show the time spent in the solver HSL MA27: The factorization time in HSL MA57 is 1.9s, compared to 1.2s for HSL MA27 (here, increasing the number of threads in the BLAS backend used by HSL MA57 does not make a difference). Hence, the block elimination algorithm implemented in HSL MA57 is not beneficial there.
When solving the KKT system, the time can be decomposed into: (1) assembling the KKT system, (2) factorizing the KKT system, and (3) computing the descent direction with triangular solves. As depicted in Figure 3, we observe that constructing the KKT system represents only a fraction of the computation time, compared to the factorization and the triangular solves. The condensed strategies LiftedKKT and HyKKT leads to faster factorization time, as the condensed KKT system can be factorized efficiently in parallel. HSL MA86 factorizes the condensed system almost 10 times faster than HSL MA27, and using cuDSS leads to an additional 4x speed-up. We focus exclusively on comparing the parallel CPU solvers (MA86 and Pardiso) for the Lifted/HyKKT formulations, as the factorization of the augmented KKT system ($K_{2}$) does not parallelize well on the CPU tasseff2019exploring . Once the KKT system is factorized, computing the descent direction with LiftedKKT is faster than with HyKKT (0.01s compared to 0.07s) as HyKKT has to run a Cg algorithm to completion to solve the Schur complement system (12) — hence requiring additional backsolves in the linear solver.
linear system linear solver build (s) factorize (s) backsolve (s) total (s)
($K_{2}$) HSL MA27 $2.85\times 10^{-2}$ $1.23\times 10^{0\phantom{-}}$ $3.48\times 10^{-1}$ $1.61\times 10^{0\phantom{-}}$
LiftedKKT (15) CHOLMOD $8.55\times 10^{-2}$ $5.84\times 10^{-1}$ $1.17\times 10^{-1}$ $7.86\times 10^{-1}$
LiftedKKT (15) pardiso $8.16\times 10^{-2}$ $2.23\times 10^{-1}$ $1.38\times 10^{-1}$ $4.43\times 10^{-1}$
LiftedKKT (15) MA86 $8.57\times 10^{-2}$ $36\times 10^{-1} 1.1$$18\times 10^{-1}$ $3.39\times 10^{-1}$
LiftedKKT (15) cuDSS $2.26\times 10^{-3}$ $4.91\times 10^{-2}$ $1.09\times 10^{-2}$ $6.23\times 10^{-2}$
HyKKT (11) CHOLMOD $8.01\times 10^{-2}$ $5.89\times 10^{-1}$ $7.22\times 10^{-1}$ $1.39\times 10^{0\phantom{-}}$
HyKKT (11) pardiso $9.06\times 10^{-2}$ $2.21\times 10^{-1}$ $9.19\times 10^{-1}$ $1.23\times 10^{0\phantom{-}}$
HyKKT (11) MA86 $8.60\times 10^{-2}$ $1.35\times 10^{-1}$ $7.50\times 10^{-1}$ $9.71\times 10^{-1}$
HyKKT (11) cuDSS $1.98\times 10^{-3}$ $3.38\times 10^{-2}$ $6.66\times 10^{-2}$ $1.02\times 10^{-1}$
Figure 3: Breakdown of the time spent in one IPM iteration for different linear solvers, when solving 78484epigrids. HSL MA27 factorizes the original augmented KKT system ($K_{2}$), and is used as a reference. The column for build shows the time spent in assembling the KKT matrix, whereas the columns for factorize and backsolve display the time spent in the numerical factorization and in the triangular solves, respectively. We set up Pardiso and HSL MA86 to run in parallel using 8 threads.
@@ 6.2 Benchmark on OPF instances
We run a benchmark on OPF instances from the PGLIB benchmark babaeinejadsarookolaee2019power . We compare the GPU solver configured with LiftedKKT and HyKKT against the CPU solver configured with the augmented KKT system and HSL MA27. For IPM solvers, we use the MadNLP solver shin2021graph with the ExaModels.jl modeling tool. We set the IPM tolerance to $10^{-6}$. Regarding HyKKT, we set $\gamma=10^{7}$, following the analysis in §6.1.2. The Dolan & Moré performance profile dolan2002benchmarking is displayed in Figure 4. The performance profile on the left compares LiftedKKT and HyKKT with MadNLP+Ma27. As a baseline, we include the performance obtained by Ipopt and Knitro, which also run with HSL MA27. Overall, the performance of HSL MA27 on the CPU is consistent with what was reported in babaeinejadsarookolaee2019power . The second performance profile on the right focuses on the performance obtained with LiftedKKT and HyKKT, using various sparse linear solvers. We observe that cuDSS leads to the fastest solution time compared to Panua Pardiso and HSL MA86 running in parallel on the CPU using 8 threads.
@@
The detailed results are displayed in Table 3: The table displays the time spent on initialization, the time spent in the linear solver, and the total solving time. On the GPU, LiftedKKT+cuDSS is faster than HyKKT+cuDSS on small and medium instances; indeed, the algorithm does not have to run a Cg algorithm at each IPM iteration, limiting the number of triangular solves performed at each iteration. Both LiftedKKT+cuDSS and HyKKT+cuDSS are significantly faster than HSL MA27. HyKKT+cuDSS is slower when solving 8387_pegase, as in this particular instance, the parameter $\gamma$ is not set high enough to reduce the total number of Cg iterations, leading to a 4x slowdown compared to LiftedKKT+cuDSS. Nevertheless, the performance of HyKKT+cuDSS improves on the largest instances, with almost an 8x speed-up compared to the reference HSL MA27.
@@
The benchmark presented in Table 3 was generated using NVIDIA A100 GPUs (80GB, current selling price: \$10k). We have also compared the performance with cheaper GPUs: a NVIDIA A1000 (a laptop-based GPU, 4GB memory) and a NVIDIA A30 (24GB memory, price: \$5k). As a comparison, the selling price of the AMD EPYC 7443 processor used for the benchmark on the CPU is \$1.2k. The results are displayed in Figure 5. We observe that the performance of the A30 and the A100 is relatively similar, and the cheaper A1000 GPU is already faster than HSL MA27 running on the CPU.
Among all nonlinear programs, the OPF problems have a very specific sparsity structure that leads to the formulation of a highly sparse KKT linear system ($K_{2}$). The super-sparse structure explains why HSL MA27 yields better results than HSL MA57, which can utilize parallelism at the BLAS level, as previously reported in tasseff2019exploring . For that reason, the numerical results obtained for the OPF instances may not be representative of the performance of HyKKT and LiftedKKT on generic nonlinear programs. In the next section, we analyze the performance of HyKKT and LiftedKKT on the CUTEst benchmark to provide a better overview of the methods’ limitations.
MadNLP+MA27 MadNLP+LiftedKKT+cuDSS MadNLP+HyKKT+cuDSS
(CPU) (GPU) (GPU)
it init AD lin total it init AD lin total it init AD lin total
1354_pegase 48 0.01 0.04 0.33 0.46 71 0.08 0.06 0.66 1.05 48 0.08 0.06 0.24 0.56
2000_goc 42 0.03 0.06 0.63 0.86 48 0.12 0.04 0.43 0.78 42 0.12 0.06 0.23 0.57
2312_goc 44 0.02 0.05 0.57 0.77 63 0.12 0.05 0.45 0.90 44 0.11 0.06 0.32 0.67
2742_goc 122 0.04 0.28 3.55 6.90 168 0.15 0.20 2.38 3.91 - - - - -
2869_pegase 55 0.05 0.11 1.00 1.42 57 0.15 0.05 0.33 0.80 55 0.14 0.08 0.40 0.88
3022_goc 55 0.03 0.10 0.92 1.26 53 0.13 0.06 0.30 0.74 55 0.13 0.07 0.37 0.81
3970_goc 48 0.06 0.14 1.89 2.64 48 0.22 0.05 0.49 1.00 48 0.20 0.09 0.40 0.92
4020_goc 59 0.06 0.18 3.56 4.46 64 0.22 0.05 0.77 1.60 59 0.21 0.12 0.67 1.33
4601_goc 71 0.06 0.22 3.08 4.11 67 0.21 0.06 0.60 1.22 71 0.21 0.14 0.67 1.40
4619_goc 49 0.07 0.18 3.07 3.95 52 0.26 0.05 0.70 1.34 49 0.25 0.10 0.54 1.16
4837_goc 59 0.06 0.20 2.44 3.39 56 0.22 0.05 0.44 1.10 59 0.22 0.11 0.57 1.21
4917_goc 63 0.06 0.19 1.85 2.75 69 0.20 0.06 0.79 1.51 63 0.20 0.10 0.54 1.16
5658_epigrids 51 0.08 0.21 2.69 3.71 48 0.27 0.05 0.55 1.23 51 0.26 0.11 0.56 1.22
7336_epigrids 50 0.11 0.26 3.44 4.64 54 0.36 0.06 1.12 2.08 50 0.37 0.11 0.58 1.36
8387_pegase 74 0.15 0.48 5.19 7.18 76 0.46 0.08 0.70 2.29 75 0.43 0.22 8.41 9.53
9241_pegase 73 0.38 0.53 5.71 8.07 107 0.50 0.12 1.18 3.42 70 0.49 0.14 1.68 2.76
9591_goc 67 0.41 0.49 10.11 12.37 75 0.49 0.09 1.99 3.74 67 0.48 0.15 1.23 2.27
10000_goc 82 0.16 0.50 5.65 7.69 64 0.39 0.07 1.12 2.21 82 0.35 0.19 1.06 2.09
10192_epigrids 54 0.42 0.46 7.18 9.42 68 0.57 0.08 1.54 3.14 54 0.51 0.14 1.08 2.09
10480_goc 71 0.40 0.60 11.01 13.68 66 0.60 0.08 1.57 3.51 71 0.56 0.14 1.61 2.74
13659_pegase 63 0.23 0.63 6.65 9.15 68 0.69 0.09 1.09 3.03 62 0.62 0.12 1.42 2.57
19402_goc 69 0.65 1.23 28.47 33.19 288 1.17 0.63 21.51 43.55 69 1.08 0.17 2.52 4.23
20758_epigrids 51 0.37 0.91 12.85 16.32 52 1.07 0.09 1.48 4.18 51 1.00 0.14 1.76 3.26
30000_goc 230 0.63 4.70 84.58 99.76 139 0.98 0.26 4.50 8.46 215 0.90 0.56 7.40 10.46
78484_epigrids 102 2.54 7.83 164.55 191.10 110 4.84 0.44 9.90 26.01 102 4.55 0.42 9.93 16.23
Table 3: OPF benchmark, solved with a tolerance tol=1e-6. The derivatives are evaluated using ExaModels.jl. The column for init displays the time spent initializing the IPM solver, including the initial symbolic factorization in the linear solver. The columns for AD, lin, and total display the time spent in the evaluation of the derivatives using AD, the time spent in the linear solver, and the total time spent in the IPM algorithm.
Refer to caption Refer to caption
@@
Figure 4: Performance profile for the PGLIB OPF benchmark, solved with a tolerance tol=1e-6. Left: comparing LiftedKKT+cuDSS and HyKKT+cuDSS with HSL MA27. Right: comparing LiftedKKT and HyKKT with various parallel sparse linear solvers.
@@
Figure 5: Comparing the performance obtained with various GPUs on three different OPF instances. We have used HyKKT to generate the results.
6.3 Benchmark on CUTEst instances
We have observed in the previous section that both LiftedKKT and HyKKT outperform state-of-the-art solvers running with HSL MA27 when running on the GPU. However, the OPF instances are specific nonlinear instances. In particular, the OPF instances exhibit extreme sparsity (with fewer than 10 non-zero coefficients per row, on average, in the augmented KKT system). Therefore, they cannot be considered representative nonlinear programs. For that reason, we complement our analysis by looking at the performance of LiftedKKT and HyKKT on large-scale CUTEst instances. Since the parallel solution approaches are intended for large-scale problems, we select the instances with more than 1,000 variables and at least 1 constraint.
As for the PGLIB benchmark, we set the IPM tolerance to $10^{-6}$ and uses $\gamma=10^{7}$ in HyKKT. We use the package CUTEst.jl to decode the SIF instances on the CPU. To use CUTEst on the GPU with HyKKT+cuDSS and Lifted+cuDSS, we have implemented a wrapper moving the sensitivities (gradient, constraints, Jacobian and Hessian) on the GPU using a copy. As a result, the evaluation of the derivatives is not as fast as if they were evaluated on the GPU, and the GPU solvers are slightly penalized by the data exchange between the CPU and the GPU.
A performance profile comparing LiftedKKT+cuDSS and HyKKT+cuDSS is provided in Figure 6. As a baseline, we include the performance obtained with Ipopt, MadNLP, and Knitro, all running with HSL MA57 (in contrast to OPF problems, HSL MA57 has better performance than HSL MA27 on CUTEst). In the performance profile (a), we observe that HyKKT and LiftedKKT are less competitive than a standard method running on the CPU, despite demonstrating robustness comparable to Knitro for LiftedKKT. However, if we select only the largest instances where Ipopt solve time is greater than 1 second (performance profile (b)), HyKKT and LiftedKKT become more competitive and are overall as fast as Knitro (the best-performing solver on the largest CUTEst instances) on 50% of the instances.
Overall, HyKKT is less robust than LiftedKKT on CUTEst: in some cases, the regularization parameter $\gamma=10^{7}$ is not set high enough to ensure a positive definite matrix $K_{\gamma}$. For instance, this is the case on ROCKET: this optimal control is known to be degenerate, in the sense that the reduced Hessian $Z^{\top}KZ$ is close to being singular (this is induced by the infamous singular arc in Goddard’s problem). We have observed that $\gamma$ must be increased to $10^{12}$ to obtain a positive definite matrix $K_{\gamma}$.
Figure 6 shows the two condensed space methods are less robust than the reference methods running on the CPU. To analyze further their performance, we select in Table 4 a dozen of representative instances where the condensed space methods are either significantly slower (unfavorable cases) or significantly faster than the references (favorable cases). Overall, the unfavorable cases correspond to one of the following situations. (i) The linear solver cuDSS takes significant time computing the symbolic factorization, as it is the case on BLOCKQP2 or ROSEPETAL2. (this explains the bump observed at the upper right of the performance profiles in Figure 6). (ii) The Jacobian has a dense column (ROCKET, BRAINPC0, BLOCKQP2, ROSEPETAL2) or almost dense columns (A5NSDSIL, A0NNDNIL, A0NSDSIL, A0NNSNSL, MPC9, CMPC10, CMPC12, DALE, CBS), leading to additional fill-in when factorizing the condensed matrices $K_{\gamma}$ and $K_{\tau}$. (iii) the equality relaxation strategy used in LiftedKKT results in a poor convergence in IPM (LUKVLE8, READING7). However, we observe on the favorable cases that if the problem does not suffer from one of the previous pitfalls, HybridKKT and LiftedKKT become competitive w.r.t. the references, with speed-up similar to what we observed before on the OPF instances (Table 3). We observe that the condensed methods are efficient to solve problems with a large number of non-zero coefficients in the Jacobian and in the Hessian (BDRY2, CLEUVEN*) or problems with a dynamical structure (MARINE, ROBOTARM). Interestingly, we observe that the condensed-space methods can efficiently accommodate dense condensed matrices $K_{\tau}$ and $K_{\gamma}$ provided they remain small enough (EIGENAU, EIGENCCO, EIGENB2, MSQRTB, where the size of the dense condensed matrix is at most $\approx 2,500\times 2,500$, compared to more $10,000\times 10,000$ in the unfavorable cases).
Refer to caption Refer to caption
(a) (b)
Figure 6: Performance profiles for the CUTEst benchmark, selecting all instances with more than 1,000 variables and at least 1 constraint. The profile (a) displays the entire benchmark, whereas the profile (b) selects only the largest instances where Ipopt takes more than 1s to run.